OpenStudy (anonymous):

am i correct? What value of k makes the equation true? 12 + (–17) = –17 + k k = 17

OpenStudy (tmak11):

so wait 12+ -17 = -17+k ...k would be 12 in that case but if not then idk

OpenStudy (mrnood):

@TheWalkingDeadFan123 rewrite the original equation but put 17 where there is k Then simplify the right hand side see if you STILL think that is correct....

OpenStudy (igreen):

\(12 + (–17) = –17 + k\) Simplify: \(12 – 17 = –17 + k\) Subtract: \(-5 = -17 + k\) Add 17 to both sides: \(k = 12\)

OpenStudy (igreen):

I don't know how you got 17, but at least you tried.

OpenStudy (igreen):

@TheWalkingDeadFan123

OpenStudy (mrnood):

@tmak11 It is discouraged in the code of conduct to give complete answers - please guide the OP but let them DO THE WORK themselves

OpenStudy (mrnood):

@iGreen ^^ the same

OpenStudy (igreen):

There is nothing against giving complete answers. The code of conduct states NOT ONLY to give out the answers. Which means you can give out the answers WITH an explanation.

OpenStudy (tmak11):

he has a point although i am sorry for not explaining

OpenStudy (mrnood):

"Give Help, Not Answers –I will encourage and guide those needing help, and not just give them an answer" "OpenStudy values the Learning process - not the ‘Give you an answer’ process •Don’t post only answers - guide the asker to a solution." Both suggest the OP is LED to a solution, not led to read the solution at the end of your explanation. The explanation is fine - but why not let OP DO the subtraction and adding 17 - it at least makes them read the working .....

OpenStudy (igreen):

I will encourage and guide those needing help, and \(NOT~JUST\) give them an answer "OpenStudy values the Learning process - not the ‘Give you an answer’ process •Don’t post \(ONLY\) answers - guide the asker to a solution."

OpenStudy (igreen):

@MrNood That might be your method of doing it, but my way is just as fine with the CoC.

OpenStudy (mrnood):

IF you give the answer at the end of your guidance many users will simply copy it and ignore th eexplanation. Let them read your guidance and USE it to do some work

OpenStudy (igreen):

@MrNood It is their responsibility to look at the explanation. I don't see that listed anywhere int he CoC, do you?

OpenStudy (igreen):

*in the

OpenStudy (mrnood):

I just strongly believe you will help the users more in the long run by letting THEM come to the answer THROUGH your guidance. "I hear and I forget. I see and I remember. I do and I understand. Confucius " However - enough said I guess

OpenStudy (anonymous):

i would agree with you @TheWalkingDeadFan123